OpenStudy (anonymous):

Help Please!!! Will give you a medal :) 1. The distance traveled by an object can be modeled by the equation d = ut + 0.5at2 where d = distance, u = initial velocity, t = time, and a = acceleration. Solve this formula for a. Show all steps in your work.

OpenStudy (anonymous):

you have to get a alone on one side of the equation. start by isolating the term with a in it (since there's only 1 term with a in it). can you do that?

OpenStudy (anonymous):

Can you walk me through it? So I would start off with d=ut+0.5at2 and I would subtract 0.5 from 2 first right?

OpenStudy (anonymous):

no... ut is on the same side as the term with a in it. you need to first get that term over to the other side.

OpenStudy (anonymous):

"that term" being ut

OpenStudy (anonymous):

Oh okay so then how would I do that?

OpenStudy (anonymous):

subtract that term from both sides

OpenStudy (anonymous):

d=ut+0.5at2 d=0.5-ut+at2? Im sorry if its getting annoying because im not understanding this.

OpenStudy (anonymous):